OpenStudy (anonymous):

Expand and simplify; -2xy(3x²y³-5x³y²) (x-3)(x²-2x+5)

OpenStudy (anonymous):

Are they two separate questions?

OpenStudy (anonymous):

\[-2xy(3x²y³-5x³y²)=-6x^{2+1}y^{3+1}+10x^{3+1}y^{2+1}\]\[=-6x^{3}y^{4}+10x^{4}y^{3}\]

OpenStudy (anonymous):

\[(x-3)(x²-2x+5)=x^3-2x^2+5x-3x^2+6x-15\]\[=x^3-5x^2+11x-15\]

OpenStudy (anonymous):

Yes they are, Thank you